OpenStudy (happykiddo):

A polyhedron has 12 faces and 30 edges. What is the number of vertices of the polyhedron? 18 20 40 42

Directrix (directrix):

Vertices + Faces = Edges + 2 (courtesty of Euler) V +12 = 30 + 2 V = 32 - 12 V = 20

OpenStudy (happykiddo):

very nice thanks for explanation too :D